OpenStudy (anonymous):

An object moving with a velocity of 25 m/s has an increase in speed of 5 m/s2 for 4 seconds. what is the object's velocity after 2 seconds? i know the answer is 35 m/s.. but can someone show me the steps on how to get there and the equations i should be using. thanks!!

OpenStudy (anonymous):

@ProfBrainstorm helpp!!!

OpenStudy (anonymous):

Initial velocity of object u='25m/s , and acceleration is constant for 4 seconds according to the data and is equal to 5m/s ^2. Since t=2 seconds lies within that time frame of 4 seconds, we use a=5 m/s^2. Since v=u +at where v is final velocity, u is initial velocity, a is acceleration and t is time, v = 25+5*2 = 35 m/s^2

OpenStudy (anonymous):

Sorry units are wrong ... It's 35m/s

OpenStudy (anonymous):

thank you so much!!!